40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 20 01:40:33 2011, Koji, Summary, SUS, MC2 Misaligned 2:27PM on Wednesday MC2_misalign.png
    Reply  Thu Oct 20 18:57:35 2011, Suresh, Summary, SUS, MC2 Misaligned 2:27PM on Wednesday 
       Reply  Fri Oct 21 02:36:44 2011, Suresh, Summary, SUS, MC2 Misaligned 2:27PM on Wednesday : MC Realigned 
       Reply  Fri Oct 21 02:57:38 2011, Suresh, Summary, SUS, MC2 Misaligned 2:27PM on Wednesday : cause traced 
          Reply  Fri Oct 21 09:08:33 2011, steve, Update, SUS, MC2 Misalignment mc2.jpg
             Reply  Wed Nov 2 10:04:05 2011, steve, Update, SUS, MC2 chamber anchoring tightened 
Message ID: 5717     Entry time: Fri Oct 21 02:36:44 2011     In reply to: 5716
Author: Suresh 
Type: Summary 
Category: SUS 
Subject: MC2 Misaligned 2:27PM on Wednesday : MC Realigned 

Quote:

Quote:

There looks some activity at around MC2 on Wednesday afternoon.
It caused the misalignment of MC2. Misalignment was not found in MC1/3.

It seems that the incident beam on the MC was aligned in the evening.
This increased the MC transmission but it is vibible that the spot on MC2 is shifted from the center.

We need an action on this issue tomorrow in the daytime.

 

I am working on fixing this.  You might some strange stuff going on in the control room screens.  Pls ignore it till I am done.

 

 

  I have realigned the MC by recentering the spots on all the MC optics.  The current spot positions (in mm) are:

MC1P     MC2P     MC3P      MC1Y      MC2Y     MC3Y

0.2245    0.3364   -0.2801   -1.8891    0.1631   -1.744

Initially the lockins 2 and 5 showed very small outputs.  This was traced to the fact that we have recently switched on a 28Hz ELP filter module in the MC2 ASC filter bank which introduces an extra phase of about 75deg..  See this elog.

When the MC ASS lockins were initially setup, the phase was set with this filter module switched off.  Since quite some time has passed since the last calibration of these phases, I readjusted the phases to minimise the  Q_OUTPUT and I also adjusted the GAINs in the SIG filter banks  of all the six lockins so that their I_OUT's drop by the calibration value of -2.65 when an offset of 0.1 is introduced into the MC suspension output matrices.  Two short scripts in the $scripts$/ASS/ directory help in setting and removing these offsets.  They are called MCxoffsetOn and MCxoffsetOff.   They have to be edited appropriately to address each DoF of the MC.

The $scripts$/ASS/mcassUp script., which sets up everything to make the MC spot decentering measurement, has been edited to set these new phases and gains.  The old settings have been commented out.

I then centered the spots on the WFS sensors and the MC_TRANS QPD.  We are now ready to make the MC WFS output matrix transfer coef measurement again, but this time with the WFS loops closed.

 

ELOG V3.1.3-